The length and width of a rectangle are in the ratio 7:3. If the perimeter is 62cm calculate the width
Answer by ankor@dixie-net.com(22740)

The length and width of a rectangle are in the ratio 7:3.
If the perimeter is 62cm calculate the width
:
let x = the multiplier
then
7x = the length
and
3x = the width
The perimeter
2(7x) + 2(3x) = 62
14x + 6x = 62
20x = 62
x = 62/20
x = 3.1 is the multiplier
:
therefore:
3.1(7) = 21.7 is the length
3.1(3) = 9.3 is the width
:
:
Check this
2(21.7) + 2(9.3) =
43.4 + 18.6 = 62